OpenStudy (anonymous):

A rectangle has a length of the cube root of 81 inches and a width of 3 to the 2 over 3 power inches. Find the area of the rectangle. 3 to the 2 over 3 power inches squared 3 to the 8 over 3 power inches squared 9 inches squared 9 to the 2 over 3 power inches squared

OpenStudy (anonymous):

@iGreen

OpenStudy (cwrw238):

width = 3^(2/3) length = cube root of 81 = 3 area = 3 * 3^(2/3) use the rule of indices:- x^n * x^m = x^(n+m)

OpenStudy (anonymous):

so it would be....

OpenStudy (cwrw238):

oops sorry 3 is not the cube root of 81 silly me!!!

OpenStudy (anonymous):

i think it is isnt it?

OpenStudy (cwrw238):

no 81 = 3*3*3*3 3 is the 4th root of 81

OpenStudy (anonymous):

so it would be 27

OpenStudy (cwrw238):

3 is cube root of 27 - yes

OpenStudy (cwrw238):

u sure you have the question right?

OpenStudy (cwrw238):

if you have have then 81 = 3^(4/3)

OpenStudy (radar):

Cube root of 81 is the cube root of (27 * 3) which is 3 * 3^(1/3) now proceed.

OpenStudy (radar):

or yes that too.

OpenStudy (cwrw238):

yes - as radar said

OpenStudy (cwrw238):

so its 3^(4/3) * 3^(2/3)

OpenStudy (anonymous):

so the answer would beeeeeeeeeeeeee...............................

OpenStudy (cwrw238):

add 2/3 + 4/3 what do you get?

OpenStudy (anonymous):

2

OpenStudy (cwrw238):

right so its 3^2

OpenStudy (cwrw238):

3^2 = ?